# Non-degenerate mixed maps and contact structures

### v1

@@ 2.3. Milnor fibrations
Oka in [26, Theorem 33] proved that strongly non-degenerate convenient mixed functions have Milnor fibrations on the tube and the sphere which are smoothly equivalent. We shall prove an analogous statement for mixed maps under the strong non-degeneracy and linear discriminant conditions. We refer the reader to [9] for further details on the subject of Milnor fibrations.
@@
We recall some definitions. A map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ admits a Milnor fibration on the tube if for each $r>0$ sufficiently small there exists $\delta=\delta(r)>0$ such that the map
(2.2) $F:\mathbb{B}_{r}\cap F^{-1}(\mathbb{S}_{\delta}^{2k-1})\longrightarrow\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}$
@@
is a locally trivial fibration over its image, where $\mathbb{B}_{r}$ is the open ball centered at the origin with radius $r$ and $\Delta_{r}$ is the discriminant of $F$ restricted to $F(\mathbb{B}_{r})$. For instance, the ICIS condition $\Sigma_{F}\cap F^{-1}(0)\subset\{0\}$ implies the existence of such a fibration by [9, Theorem 2.3]. Hence, the maps in Theorem 2.10 admit a fibration on the tube.
In the notation above, the map $F$ admits a fibration on the sphere if for each sufficiently small $r>0$ there exists a $\delta=\delta(r)>0$ such that the map
(2.3) $\frac{F}{\|F\|}:\mathbb{S}_{r}^{2n-1}\cap F^{-1}(\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r})\longrightarrow\mathbb{S}^{2k-1}_{\delta}\setminus\Delta_{r}$
is a locally trivial fibration over its image.
Lemma 2.13.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a strongly non-degenerate mixed map germ such that for all non-empty $I\subset\{1,\dots,n\}$ the following conditions hold:
@@
    If $|I|<k$, then $\mathbb{C}^{I}\subset\Sigma_{F}$.
@@
    If $|I|\geq k$, the coordinate functions satisfy $f^{i,I}\not\equiv 0$.
Then there exists $r_{0}>0$ such that the fibers of the map
$F:\mathbb{B}_{r}\cap F^{-1}\left(\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}\right)\longrightarrow\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}$
are transversal to the spheres $\mathbb{S}^{2n-1}_{r}$ for all $0<r<r_{0}$ and $\delta=\delta(r)>0$ sufficiently small.
@@
If the statement is false, by [31, Proposition 7] and the Curve Selection Lemma, there exists real analytic curves $w(t)\subset\mathbb{C}^{n}\setminus F^{-1}(\Delta_{F})$ and $\beta_{1}(t),\dots,\beta_{k}(t)\in\mathbb{C}$, $\lambda(t)\subset\mathbb{R}$ non-simultaneously vanishing such that:
$\lambda(t)w(t)=\sum_{j=1}^{k}\beta_{j}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{j}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{j}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)).$
Let us denote $I=\{j:w_{j}(t)\not\equiv 0\}$, where $|I|\geq k$, since $w(t)\cap\Sigma_{F}=\emptyset$. This also implies $\lambda(t)\not\equiv 0$. Under these conditions, $f^{i,I}\not\equiv 0$ and we may write the analytic expansions of the coordinate functions and curves:
$f^{i,I}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ $=c_{i}t^{r_{i}}+o(t),\;c_{i}\in\mathbb{C}^{*}\,,r_{i}\geq 1,$
$\lambda(t)$ $=\lambda_{0}t^{s}+o(t),\;\lambda_{0}\in\mathbb{R}^{*}\,,s\geq 1,$
@@
$\beta_{i}(t)$ $=b_{i}t^{q_{i}}+o(t),\;b_{i}\in\mathbb{C}^{*}\,,q_{i}\geq 1,$
for all $i\in I$. We may suppose $q_{1}+d_{1}\leq\dots\leq q_{k}+d_{k}$, where $d_{i}$ is the weighted homogeneous degree of $f^{i,I}_{P}$. For each $i$, the relation above becomes
@@
$a_{i}\lambda_{0}t^{s+p_{i}}+o(t)=\left(b_{1}\frac{\partial f_{P}^{1,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{1}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{1,I}}{\partial z_{i}}}\right)t^{d_{1}+q_{1}-p_{i}}+o(t)+\dots+\left(b_{k}\frac{\partial f_{P}^{k,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{k}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{k,I}}{\partial z_{i}}}\right)t^{d_{k}+q_{k}-p_{i}}+o(t).$
It follows that:
@@
$\sum_{j=1}^{l}b_{j}\frac{\partial f^{j,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{j,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\begin{cases}0,&\,\text{if}\;q_{1}+d_{1}-p_{i}-r_{i}<s+p_{i}\\ \lambda_{0}a_{i},&\,\text{if}\;q_{1}+d_{1}-p_{i}-r_{i}=s+p_{i}\end{cases}$
Let us define $K=\{i:q_{1}+d_{1}-p_{i}-r_{i}=s+p_{i}\}$ and we claim that $K\neq\emptyset$. On the contrary, $a\in\mathbb{C}^{*m}$ becomes a critical point of $F_{P}^{I}$ and we apply the first item of Proposition 2.9 to get the first contradiction. The second follows from the following argument. Consider the equations:
@@
(2.4) $\left\langle\sum_{i=1}^{k}\beta_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}F(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muDF}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)),w^{\prime}(t)\right\rangle=\frac{1}{2}\lambda(t)\frac{d}{dt}\|w(t)\|^{2}.$
We may split this dot product as follows:
(2.5) $\left\langle\sum_{i=1}^{k}{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{i},\beta_{i}\frac{dw}{dt}\right\rangle$ $=\sum_{i=1}^{k}\left\langle{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{i}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),b_{i}P\cdot a\right\rangle t^{d_{i}-1+q_{i}}+o(t),$
@@
(2.6) $\left\langle\sum_{i=1}^{k}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{i}},{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{i}\frac{d{}\mkern 3.0mu\overline{\mkern-3.0muw}}{dt}\right\rangle$ $=\sum_{i=1}^{k}\left\langle{}\mkern 3.0mu\overline{\mkern-3.0muDf^{i}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}P\cdot{}\mkern 3.0mu\overline{\mkern-3.0mua}\right\rangle t^{d_{i}-1+q_{i}}+o(t),$
where $P=(p_{1},\dots,p_{n})$ and $a=(a_{1},\dots,a_{n})$. Notice that the Hermitian product satisfies
$\operatorname{Re}\langle v,w\rangle+\operatorname{Re}\langle{}\mkern 3.0mu\overline{\mkern-3.0muv},u\rangle=\operatorname{Re}\langle v,w-{}\mkern 3.0mu\overline{\mkern-3.0muu}\rangle\;\forall\;v,w,u\in\mathbb{C}^{n}.$
Thus, we sum both products (2.5) and (2.6) to obtain that the real part of the coefficient of the lowest degree term of (2.4) is $\lambda_{0}\sum_{j\in K}\|a_{j}\|^{2}p_{j}\neq 0$, provided $K\neq\emptyset$. This allows us to compare the orders:
$d_{1}-1+q_{1}=2p+s-1,$
where $p=\min_{i}\{p_{i}\}$. On the other hand, $s+p=q_{1}+d_{1}-p-r_{l}$ for some index $l\in K$. This leads to $r_{l}=0$, which is a contradiction. ∎
Example 2.14.
Let $f^{1},\dots,f^{k}$ be mixed function germs such that $f^{j}:(\mathbb{C}^{n_{j}},0)\longrightarrow(\mathbb{C},0)$ is strongly non-degenerate, has a critical point at the origin, and $n_{j}>k$ for every $j=1,\dots,k$. Define the mixed map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n_{1}}\times\dots\times\mathbb{C}^{n_{k}},0)\longrightarrow(\mathbb{C}^{k},0)$ on separable variables. We have seen in Example 2.14 that $F$ is strongly non-degenerate. In addition, the matrix $\left(DF\right)^{I}$ does not have maximal rank if $|I|<k$. Moreover, $f^{j,I}\not\equiv 0$ for all $I$ such that $|I|\geq k$. Thus $F$ satisfies the conditions of Lemma 2.13.
@@
Let $F:\mathbb{C}^{n}\longrightarrow\mathbb{C}^{k}$ be a strongly non-degenerate mixed map as in Lemma 2.13. Suppose further that $F$ has a linear discriminant. Then $F$ admits Milnor fibrations on the tube and the sphere, and these are smoothly equivalent.
Proof.
Lemma 2.13 implies that the map (2.2) is a submersion and the existence of a tube fibration is a consequence of Ehresmann fibration theorem. Moreover, the transversality of the fibers with every small sphere yields that the map (2.3) is a fibration by [9, Proposition 2.12]. Lastly, the equivalence is established in [9, Theorem 2.16]. ∎

### v3

@@ 2.3. Milnor sets
Oka in [30, Theorem 33] proved that strongly non-degenerate convenient mixed functions have Milnor fibrations on the tube and the sphere, which are smoothly equivalent. We shall relate the non-degeneracy property to the Milnor sets of a mixed map germ, and then derive the existence of a tube fibration under certain conditions. We refer the reader to [10] and [2] for further details on the subject of Milnor fibrations.
@@
We recall some definitions. A map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ admits a Milnor fibration on the tube if, for each $r>0$ sufficiently small, there exists $\delta=\delta(r)>0$ such that the map
(2.5) $F:\mathbb{B}_{r}\cap F^{-1}(\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r})\longrightarrow\mathbb{S}_{\delta}^{2k-1}\setminus\Delta_{r}$
@@
is a locally trivial fibration over its image, where $\mathbb{B}_{r}$ is the closed ball centered at the origin with radius $r$ and $\Delta_{r}=F(\Sigma_{F}\cap\mathbb{B}_{r})$ is the discriminant of $F$ restricted to $\mathbb{B}_{r}$. The existence of Milnor fibrations is related to the following set.
Let $\rho(z)=\sum_{i=1}^{n}\|z\|^{2}$ be the square distance function and consider an open neighborhood $\mathring{\mathbb{B}}_{r}\subset\mathbb{C}^{n}$ of the origin. Consider the following set:
$M(F):=\{z\in\mathring{\mathbb{B}}_{r}:F\not\pitchfork_{z}\rho\}.$
Notice that $M(F)$ consists of the critical points of the map $(F,\rho):\mathring{\mathbb{B}}_{r}\longrightarrow\mathbb{C}^{k}\times\mathbb{R}$. Hence, this set is closely related to the transversality property. In general, $M(F)$ is not well-defined as a set germ (see, for instance, [2]). In the affirmative case, it is called the Milnor set of $F$. For instance, if the image $\text{Img}(F)$ of $F$ and the discriminant $\Delta_{F}:=F(\Sigma_{F})$ are well-defined as set germs, a condition called nice, the Milnor set is also well-defined. For details, we refer the reader to [2], [18], and [35], and the references therein. By [2, Lemma 3.3], the following condition ensures the existence of the Milnor fibration on the tube through the Milnor set:
(2.6) $\overline{M(F)\setminus F^{-1}(\Delta_{F})}\cap V_{F}\subseteq\{0\}.$
We discuss the Milnor fibration on the sphere. Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a nice mixed map germ and suppose that $\Delta_{F}$ is linear, which means that it is a union of lines passing through the origin. This is also known as radial discriminant. Let $\mathring{\mathbb{B}}_{r}\subset\mathbb{C}^{n}$ be an open neighborhood of the origin and consider the map:
(2.7) $\phi_{F}:=\frac{F}{\|F\|}:\mathring{\mathbb{B}}_{r}\setminus F^{-1}(\Delta_{F})\longrightarrow\mathbb{S}^{2k-1}.$
We say that $F$ admits a Milnor fibration on the sphere if, for every sufficiently small $r>0$, its restriction
(2.8) $\frac{F}{\|F\|}:\mathbb{S}_{r}^{2n-1}\setminus F^{-1}(\Delta_{F})\longrightarrow\mathbb{S}^{k-1}\setminus\pi_{\eta}(\Delta_{F})$
is a locally trivial fibration, where $\pi_{\eta}:\mathbb{S}_{\eta}^{2k-1}\longrightarrow\mathbb{S}^{k-1}$ is the normalization map and $\eta>0$ is a linearity radius. In [10, Theorems 2.13 and 2.16] and [3] is shown the following. If $F$ admits a Milnor fibration on the tube, and in addition, the map (2.8) is a submersion for every sufficiently small $r>0$, then there exists a Milnor fibration on the sphere. This condition is equivalent to the Milnor set $M(\phi_{F})\subset F^{-1}(\Delta_{F})$ and is called $d$-regularity in [10] and $\rho$-regularity in [2]. Moreover, these fibrations are smoothly equivalent, provided that we restrict (2.5) to the interior $\mathring{\mathbb{B}}_{r}$ of the ball $\mathbb{B}_{r}$. The most common examples of maps for which (2.8) is a submersion are those with the radial weighted homogeneous property (see [3] for details).
Remark 2.26.
Assume that $F$ is a nice mixed map germ and let $S=M(F),M(\phi_{F})$. Observe that $\Sigma_{F}\subset F^{-1}(\Delta_{F})\cap S$. Moreover, suppose that $0\not\in\overline{S\setminus F^{-1}(\Delta_{F})}$. Then there exists $r>0$ such that $\mathring{\mathbb{B}}_{r}\cap S=\mathring{\mathbb{B}}_{r}\cap F^{-1}(\Delta_{F})$, where $\mathring{\mathbb{B}}_{r}\subset\mathbb{C}^{n}$ is an open neighborhood of the origin. Hence, we obtain that $S\setminus F^{-1}(\Delta_{F})=\emptyset$ as a set germ at the origin.
We recall the following notation. Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ and $I\subset\{1,\dots,n\}$ a non-empty subset. Then $J_{I}=\{j:f^{j,I}\not\equiv 0\}$ and $\left(J_{I}\right)^{C}$ denotes its complement. Moreover, the set of non-vanishing subspaces of $F$ is defined by:
$\mathcal{I}_{F}=\{I\subset\{1,\dots,n\}:\left(J_{I}\right)^{C}=\emptyset\}.$
Definition 2.12 and the notion of non-vanishing spaces motivate the following.
Definition 2.27.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a nice mixed map germ. We say that $F$ is Milnor convenient with respect to $S$, where $S$ denotes $M(F)$ or $M(\phi_{F})$, if the following condition hold for every $I\notin\mathcal{I}_{F}$:
$\overline{S\setminus F^{-1}(\Delta_{F})}\cap V_{F}\cap\mathbb{C}^{I}\subseteq\{0\}.$
For instance, if $F$ is convenient as in Definition 2.12, then $F$ is also Milnor convenient with respect to $M(F)$ and $M(\phi_{F})$. In [3, Example 6 and Lemma 6] are presented examples of nice mixed map germs $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n-1},0)$ such that $V_{F}\cap\mathbb{C}^{I}\neq\{0\}$ for $|I|=k$, but the condition above holds for $S=M(F)$. The main result of this subsection relates the Newton non-degeneracy property and Milnor sets.
Theorem 2.28.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate nice mixed map germ and let $S$ denote $M(F)$ or $M(\phi_{F})$.
@@
    If $F$ is convenient, then $S\cap V_{F}=\{0\}$.
@@
    If $F$ is Milnor convenient with respect to $S$, then $\overline{S\setminus F^{-1}(\Delta_{F}})\cap V_{F}\subseteq\{0\}$.
In particular, if $F$ is Milnor convenient with respect to $M(F)$, then it admits a Milnor fibration on the tube.
@@
We present the proof of the second item, which also applies to the first one. Let $S=M(F)$ or $S=M(\phi_{F})$. We may suppose $\overline{S\setminus F^{-1}(\Delta_{F}})\neq\emptyset$. The argument is based on the proofs of [30, Lemma 31] and [9, Theorem 3.5]. If this statement is false, by the characterization of points in the Milnor sets presented in [34, Propositions 5 and 7] and the Curve Selection Lemma, there exists a real analytic curve $w(t)\subset\overline{S\setminus F^{-1}(\Delta_{F}})\cap V_{F}$ such that $\lim_{t\to 0}w(t)=0$ and others real analytic curves $\beta_{1}(t),\dots,\beta_{k}(t)\in\mathbb{C},\lambda(t)\subset$ non-simultaneously vanishing such that:
(2.9) $\lambda(t)w(t)=\sum_{j=1}^{k}\beta_{j}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{j}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{j}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{j}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)).$
Let us denote $I=\{i:w_{i}(t)\not\equiv 0\}$. We may write the analytic expansions of the curves:
$\lambda(t)$ $=\lambda_{0}t^{s}+o(t),\;\lambda_{0}\in\mathbb{R}\,,s\geq 1,$
@@
$\beta_{j}(t)$ $=b_{j}t^{q_{j}}+o(t),\;b_{j}\in\mathbb{C}^{*}\,,q_{j}\geq 1,$
for all $i\in I$ and $j$ such that $\beta_{j}\not\equiv 0$. For each $i$, the relation above becomes
@@
$a_{i}\lambda_{0}t^{s+p_{i}}+o(t)=\left(b_{1}\frac{\partial f_{P}^{1,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{1}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{1,I}}{\partial z_{i}}}\right)t^{d_{1}+q_{1}-p_{i}}+o(t)+\dots+\left(b_{k}\frac{\partial f_{P}^{k,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{k}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f_{P}^{k,I}}{\partial z_{i}}}\right)t^{d_{k}+q_{k}-p_{i}}+o(t),$
where $d_{j}$ is the weighted homogeneous degree of $f^{j,I}_{P}$. We may suppose $q_{1}+d_{1}\leq\dots\leq q_{k}+d_{k}$. Moreover, let $l\leq k$ such that $q_{1}+d_{1}=\dots=q_{l}+d_{l}$. Without loss of generality, suppose that $\beta_{j}\not\equiv 0$ for every $j=1,\dots,l$. It follows that:
@@
(2.10) $\sum_{j=1}^{l}b_{j}\frac{\partial f^{j,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{j,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})=\begin{cases}0,&\,\text{if}\;\lambda(t)\equiv 0,\\ 0,&\,\text{if}\;q_{1}+d_{1}-p_{i}<s+p_{i},\\ \lambda_{0}a_{i},&\,\text{if}\;q_{1}+d_{1}-p_{i}=s+p_{i}.\end{cases}$
Let us define $K=\{i:q_{1}+d_{1}-p_{i}=s+p_{i}\}$. We claim that $\lambda(t)\not\equiv 0$ and $K\neq\emptyset$. In this case, $w(t)\subset V_{F}$, and then $a\in\mathbb{C}^{*I}\cap\left(F^{I}\right)_{P}^{-1}(0)$ is a critical point of $\left(F^{I}\right)_{P}$. By the Milnor convenience, we obtain that $F^{I}$ is non-degenerate, which is a contradiction. Consider now the following expression:
@@
(2.11) $\left\langle\sum_{i=1}^{k}\beta_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}F(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))+{}\mkern 3.0mu\overline{\mkern-3.0mu\beta}_{i}(t){}\mkern 3.0mu\overline{\mkern-3.0muDF}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)),w^{\prime}(t)\right\rangle=\lambda(t)\langle w(t),w^{\prime}(t)\rangle.$
The lower degree term on the left-hand side of (2.11) is $q_{1}+d_{1}-1=\dots=q_{l}+d_{l}-1$, and its coefficients are
$\sum_{i=1}^{n}\sum_{j=1}^{l}\left(b_{j}\frac{\partial f^{j,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{j,I}}{\partial z_{i}}}\right){}\mkern 3.0mu\overline{\mkern-3.0mua}_{i}p_{i}.$
By (2.10), this sum is equal to
(2.12) $\sum_{i\in K}\sum_{j=1}^{l}\left(b_{j}\frac{\partial f^{j,I}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}+{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{j,I}}{\partial z_{i}}}\right){}\mkern 3.0mu\overline{\mkern-3.0mua}_{i}p_{i}=\lambda_{0}\sum_{i\in K}\|a_{i}\|^{2}p_{i}\neq 0,$
since $K\neq\emptyset$. We may rewrite it as
$\sum_{j=1}^{l}\langle b_{j}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{j,I}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle+\langle{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j,I}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle\neq 0,$
where $Pa=(p_{1}a_{1},\dots,p_{n}a_{n})$. Hence, (2.12) implies that
@@
(2.13) $\operatorname{Re}\left(\langle b_{j_{0}}{}\mkern 3.0mu\overline{\mkern-3.0muD}f_{P}^{j_{0},I}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle+\langle{}\mkern 3.0mu\overline{\mkern-3.0mub}_{j_{0}}{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j_{0},I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle\right)\neq 0$
for some $j_{0}$. On the other hand, since $w(t)\subset V_{F}$, one has $f^{j_{0}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\equiv 0$, and thus
(2.14) $\frac{d}{dt}f^{j_{0},I}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))=\left(\langle{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j_{0},I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle+\langle{}\mkern 3.0mu\overline{\mkern-3.0mu{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{j_{0},I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),P{}\mkern 3.0mu\overline{\mkern-3.0mua}\rangle\right)t^{d_{j_{0}}-1}+o(t)\equiv 0.$
Then, we obtain that:
(2.15) $\langle{}\mkern 3.0mu\overline{\mkern-3.0muDf^{j_{0},I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),Pa\rangle+\langle{}\mkern 3.0mu\overline{\mkern-3.0mu{}\mkern 3.0mu\overline{\mkern-3.0muD}f^{j_{0},I}_{P}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua}),P{}\mkern 3.0mu\overline{\mkern-3.0mua}\rangle=0.$
If we multiply (2.15) by $\mkern 3.0mu\overline{\mkern-3.0mub}_{j_{0}}$, the resulting equation has real part equal to (2.13), which leads to a contradiction, and we conclude the proof. ∎
An alternative statement for Theorem 2.28, which has the same proof and does not have the Milnor convenience as an assumption, is the following.
@@
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate nice mixed map germ and let $S$ denote $M(F)$ or $M(\phi_{F})$. Then $S\cap V_{F}\cap\mathbb{C}^{*I}=\emptyset$ for every $I\in\mathcal{I}_{F}$. In particular, if $F$ is convenient, then $S\cap V_{F}=\{0\}$ and $F$ admits a Milnor fibration on the tube.
Since $\Sigma_{F}\subset M(F)$, the previous result implies Theorem 2.13 under the convenience hypothesis. Moreover, by the criterion (2.6), this leads to the existence of Milnor fibrations on the tube for nondegenerate and convenient mixed maps. Actually, the property $\Sigma_{F}\cap V_{F}=\{0\}$, ensured by Theorem 2.13, already leads to this conclusion, by [10, Theorem 2.3]. In virtue of the radial weighted homogeneous property of mixed Pham-Brieskorn polynomials, we can derive the following statement.
Corollary 2.30.
A mixed Hamm map germ admits a Milnor fibration on the tube and the sphere. Moreover, these fibrations are smoothly equivalent, once we restrict the tube to the interior of a sufficiently small open ball around the origin.